# Odd-One-Out: Anomaly Detection by Comparing with Neighbors

### v1

@@ 4.1 Datasets
As no prior dataset exists for the task, we propose two challenging scene datasets, both essential for thorough evaluation: ToysAD-8K and PartsAD-15K. ToysAD-8K includes real-world objects from multiple categories. This allows us to evaluate our model’s ability to generalize to unseen object categories. PartsAD-15K comprises a more diverse collection of mechanical object parts with arbitrary shapes, thus being free from any class-level inductive biases. Both datasets include a wide range of fine-grained anomaly instances motivated by real-world applications in inspection and quality control. Scenes are generated with diverse backgrounds, illuminations, and camera viewpoints using photo-realistic ray tracing [43]. Next, we discuss the data generation for both datasets.
@@
ToysAD-8K. We begin with a subset of $1050$ shapes from the Toys4K dataset [41]. The subset covers a wide range of objects from $51$ categories. We automatically create anomalies for a given 3D shape by applying various deformations to both the geometry and texture. This includes generating realistic cracks and fractures using [39], applying random geometric deformations [43] like bumps, bends, and twists, as well as randomly translating, rotating, and swapping materials in different parts of the shapes. In total, we generated $2345$ anomaly shapes. To generate each scene, we first randomly choose a set of objects consisting of both normal and their anomaly of the same instances. We ensure that the majority of objects in each scene are normal. We also have few scenes where no anomaly is present. The rotational poses for the objects are obtained using rigid body simulation [4]. The objects are scaled and placed into the scene at random locations, ensuring collisions do not occur. We generated total $8K$ scenes. Each scene consists of $3$-$6$ objects rendered in $20$ views. For the training set, we randomly select $5K$ scenes from the $39$ categories. We build two disjoint test sets. The first one (seen) contains $1K$ scenes from the seen categories but with unseen object instances. The second one (unseen) contains $2K$ scenes from rest of the $12$ novel categories.
@@
PartsAD-15K. We use a subset of the ABC dataset [27] that consists of $4200$ shapes. We follow the strategy above to generate anomalies. Additionally, for each shape, we sample geometrically close instances from the dataset and assign them as anomalies to use in the same scene. This approach generates a large set of high-quality anomalies that closely resemble their normal counterparts in high-level geometry, but with subtle shape variations that make them anomalous in the context of the normal ones. In total, we generated $10,203$ anomaly shapes. Using these shapes, we created $15K$ scenes, each consisting of $3$ to $12$ objects rendered from $20$ different viewpoints. We divide the dataset into a $12K$ training set and the rest as the test set. Note that the proposed datasets, source code, and models will be made public based upon publication.
4.2 Implementation details
We use a ResNet50-FPN [30] as our 2D encoder backbone. Our 3D backbone consists of a four-scale encoder-decoder-based 3D CNN [32]. We use $M=5$ images as input, each with a resolution of $256\times 256$. Our architecture is flexible to accept a different $M$ during inference. During training, we consider a total of $2M$ views, which we separate into two sets, each containing $M$ views. We use one set to build the neural volume and the cameras of the other set to render the results, and vice versa. The 3D volume $\bm{F}_{v}$ contains $96\times 96\times 16$ voxels with a voxel size of $4cm$. We sample $128$ points on each ray for rendering. We render the features with a spatial dimension of $32\times 32$. The $I_{t\sigma}$ in Eq. 3 is the ground truth segmentation mask of the input image, we only use it for training. The threshold applied to the density volume $\bm{V}_{\sigma}$ is set to be $0.2$ and we run the DBScan algorithm with its default parameters. We resize the teacher’s (DINOv2) features to the same spatial dimension for loss computation. These features are pre-computed for all scenes using publicly available weights, which are not updated during distillation. These DINOv2 features are then reduced to $d_{f}$=$128$ channel dimensions using PCA before distillation. We employ three sparse voxel attention blocks, and each applies 8-headed attention. The value of $k$ is chosen as $20$. The loss weights $\lambda_{\sigma}$, $\lambda_{f}$, and $\lambda_{p}$ are all set to $1$. We first pretrain the network with only the reconstruction loss for $50$ epochs. Then, we train the network end-to-end with both the reconstruction loss and the binary classification loss for another $50$ epochs. We maintain a batch size of $4$, and use the Adam optimizer with a learning rate $2\times 10^{-5}$. Finally, the run-time of our method is $65$ms on a single A40 GPU for a typical scene with 5 views as input.
4.3 Baseline Comparisons
We quantitatively evaluate the anomaly classification results using two evaluation metrics – the area under the ROC curve (AUC) and accuracy. A prediction is considered correct if the bounding box IoU is greater than 0.5 and the corresponding anomaly classification is correct. We have not designed a separate localization metric as our estimated bounding boxes are very accurate. These metrics are calculated object-wise and then averaged across all test scenes. Since no prior work exists on this task, we define several competitive baselines for comparison. Specifically, in  Tab. 1, we compare our method with two relevant approaches: a reconstruction-based baseline and two multi-view 3D object detection methods. For the reconstruction-based design, we first employ COLMAP [38] to obtain a point cloud reconstruction of the multi-object scene environment. We use default dense reconstruction parameters but utilize the provided ground truth camera matrices. Since 5 views are insufficient, we use a total of 20 views to reconstruct the scene. We extract individual object point clouds from the reconstructed scenes, and train a Siamese style network to obtain their pairwise similarity. We use DGCNN [49] (pretrained on ShapeNet [14]) as the point cloud feature extractor and the triplet loss [22] to supervise the network. Finally, we employ a voting strategy to aggregate the pairwise distances of all objects in the scene and obtain their individual binary labels. We observe that the accuracy of this method is sensitive to the reconstruction quality, leading to poor performance on both datasets. Next, we adapt ImVoxelNet [37] and DETR3D [48], two multi-view object detection frameworks in our problem setting, aiming to locate and classify each object in the scene as either anomaly or normal. ImVoxelNet uses the similar 2D-3D projection as ours to construct the voxel representation and then a 3D detection head [29] outputs the final prediction. DETR3D is a transformer-based design and uses the set prediction loss [9] for end-to-end detection without NMS. We use the ground truth 3D bounding boxes to train these models and the reported scores are calculated object-wise based on the output of the classification head. We observe that while both methods perform well for large cracks or fractures, they struggle when intra-group comparison is necessary. This is because they tend to memorize certain anomaly types without learning to generalize to compare with other objects in the scene. Our method significantly outperforms all baselines (Tab. 1) on both datasets, highlighting the effectiveness of our dedicated architecture for matching corresponding regions. Notably, the performance drop on the unseen set is relatively lower in our case, attributed to the robust 3D representation that effectively generalizes to novel categories. We show qualitative results of our method in Fig. 3
@@ 4.3 Baseline Comparisons
Datasets COLMAP [38] ImVoxelNet [37] DETR3D [48] Ours
@@
Figure 3: AD qualitative results on (a) the unseen test categories of ToysAD-8K and (b) the test set of PartsAD-15K using our proposed framework. The green box denotes correct predictions, while the red box indicates where the model misses the anomaly. Due to limited space, one view is shown per scene. Please refer to the supplementary materials for more results.
@@ 4.4 Ablations and Model Analysis
We conduct several studies into the performance of the proposed model including changes to architecture, robustness analysis and real-world experiments. Unless stated otherwise, all experiments in this section are carried out on the ToysAD-8K unseen set using 5 input views.
Figure 4: Correspondences are obtained in 3D space using the neural feature field $\bm{V}_{f}$, then projected onto 2D views using camera matrices for visualization. The feature field is rendered at respective viewpoints, the first 3 PCA components are used, with each matched to a different color channel.
Ablation of architecture design. We ablate the core components in our model and report the results in Tab. 2. All variants include the 3D feature fusion module and are optimized at least for the image reconstruction loss, which is essential for constructing scene geometry. Variant A directly maps object-centric features to their binary labels using an MLP without comparing them, while for variant B, standard attention layers are applied to the object-centric features to learn cross-object correlations. Despite the attention layers, B only shows a slight improvement (+2.1% AUC) over A.
Table 2: Ablation Results on ToysAD-8K
@@
We then introduce DINOv2 feature distillation (variant C), which significantly boosts performance by 5.8% AUC and 11.3% accuracy, indicating the importance of the part-aware 3D representation and correspondences for our task (also see Fig. 4). Our final design achieves the performance gain by utilizing sparse voxel attention, which focuses on the top-$k$ most relevant features. This leverages robust correspondences learned through DINOv2 feature distillation, effectively eliminating noisy correlations and directing attention solely to corresponding object regions.
@@
Robustness. Our method achieves some robustness to occlusion by effectively using input from multiple viewpoints. Fig. 7 illustrates examples where our model accurately identifies the anomalous region, even when occluded in some views. In Fig. 5 (left), we show how well our model can perform with additional or reduced input views at test time. We train our model with 5 views and then evaluate it using 1, 3, 5, 10 and 20 views. These results clearly indicate that our model can perform reasonably well with only 3-5 views; however, additional views can be used to boost performance at test time.
Figure 5: Impact of number of views (left) and object count (right) on model performance.
We analyze the impact of object count on model performance (see Fig. 5 (right)) using the PartsAD-15K dataset. Increasing the number of objects in a scene usually leads to more occlusion and lower resolution for each individual object. However, the performance drop between the two extremes is minimal ($<$ 1.5% AUC) as shown in the figure. In addition, we investigate the model’s ability to generalize to more objects at test time than it observed during training. To this end, we trained our model on scenes with 3-7 objects and tested on two sets: one with 3-7 objects (AUC: 86.75) and another with 8-12 objects (AUC: 85.10). These results demonstrate our model’s adaptability to varying object counts.
In another study, we test our primary objective of choosing the majority group as normal and the rest as anomalies using an example shown in Fig. 6. We create five scenes (a-e) by using two geometrically similar objects, and gradually introduce more from one (1 to 5 respectively) while maintaining a total of six objects. For example, in scene (a), the first object appears only once, making it an anomaly. Similarly, in scene (e), the second object appears only once, hence considered an anomaly. We utilize a consistent background across all scenes to ensure uniformity. As shown in the figure, our model correctly classifies the anomalies in each scene. We note scene (c) presents an ambiguous case, where both objects appear in equal numbers. Despite this, our model is able to separate the two groups.
Real world testing. Here we apply our model, which is trained on the synthetic dataset, on a small set of real test scenes. Each scene is set up in an indoor environment with adequate lighting and is captured using a 3D scanning software [1]. This results in a set of input views with globally optimized cameras. Fig. 8 illustrates the results for three such scenes.
Limitations. Our benchmark and model also have a few limitations. Firstly, this work focuses solely on a limited set of anomalies that are common to manufacturing scenarios; potentially missing other anomalies in real-world scenarios. Since acquiring real damaged objects is expensive and difficult, our dataset primarily uses shapes of synthetic objects. Moreover, our model assumes that object instances are rigid and cannot handle articulations or deformations. It also assumes that objects are not touching or completely occluded in space. Moreover, a scene with mostly anomalies can be challenging for our model due to the lack of ‘normal’ data points for comparison (except for fractures or cracks). Moreover, the performance is dependent on the anomalous region being captured in at least one view. Finally, in real-world scene testing, the noisy camera poses as well as the gap between synthetic and real-world environments may degrade the performance.
Figure 6: We create five scenes (a-e) with two similar objects, varying their counts per scene while maintaining a constant total. Our model correctly identifies minorities as anomalies in all cases except (c), where the equal number of objects creates ambiguity. Our model still selects one group.
@@
Figure 7: Resolving (a) occlusion and (b) 3D ambiguity using multi-view images. The anomaly ‘sheep’ in (a) has a missing tail (only visible in the last view due to occlusion), and the ‘hammer’ handle in (b) is bent (only apparent from the last view-angle due to 3D ambiguities).
@@
Figure 8: Three real-world scenes are tested using our method and it can successfully detect all anomalous instances. Five input views are used for testing, while one view for each scene is used for visualization.
@@ 5 Conclusion
In this paper, we have introduced a novel AD problem inspired by real-world applications along with two new benchmarks. The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple camera viewpoints to identify the ‘odd-looking’ minority group. We show that our model is robust to varying number of views and objects, and outperforms the baselines that do not consider cross-object correlations.
Broader Impacts.
The proposed techniques could be potentially used in improving quality checks and product safety by providing warnings to the users in production lines. The authors are not aware of any potential harm that may arise when the technology is used.
Appendix A Appendix / supplemental material
This appendix is structured as follows: we present additional details of the proposed datasets in Sec. A.1, training details in Sec. A.2, and additional qualitative results in Sec. A.3.
A.1 Data Generation Details
Our proposed scene AD datasets, ToysAD-8K and PartsAD-15K, are built upon two publicly available 3D shape datasets: Toys4K [41] (Creative Commons and royalty-free licenses) and ABC [27] (MIT license). For the ToysAD-8K, we selected 1,050 shapes from the Toys4K dataset, focusing on the most common real-world objects across 51 categories. A complete list of these categories is provided in Table 3. On the other hand, PartsAD-15K is a non-categorical dataset. For this dataset, we randomly selected a subset of 4,200 shapes from the large-scale ABC.
For both datasets, we consider the following anomaly types: cracks, fractures, geometric deformations (e.g., bumps, bends, and twists), translation, rotation, material mismatch, and missing parts. Additionally, for PartsAD-15K, we use a different but geometrically similar instance in the same scene as an anomaly. To obtain a geometrically similar shape, we use feature-based KNN clustering. Specifically, we extract DINOv2 features from multi-view images (rendered from multiple fixed viewpoints) of the 3D shapes, concatenate these features, and use the resulting concatenated features to build the KNN cluster. Then, we retrieve a similar 3D shape by querying the KNN cluster. To ensure the retrieved shape is geometrically similar, we calculate the Chamfer distance between the shapes and only accept a shape if the distance is below a certain threshold.
Table 3: Dataset composition of ToysAD-8K
Categories
Seen dinosaur, fish, frog, monkey, light, lizard, orange, boat, dog, lion, pig, cookie, panda, chicken,
orange, ice, horse, car, airplane, cake, shark, donut, hat, cow, apple, bowl, hamburger, octopus,
giraffe, chess, bread, butterfly, cupcake, bunny, elephant, fox, deer, bus, bottle
Unseen mug, plate, robot, glass, sheep, shoe, train, banana, cup, key, penguin, hammer
Our anomaly generation process is automatic. To ensure the quality of the generated anomalies, we perform several checks. For example, during positional or rotational anomaly creation, if a part detaches from the main body during deformation, we reject the anomaly and try again with adjusted parameters. Similarly, if removing a part makes the shape impractical, we discard it and try removing a different part. For fracture anomaly, if a particular fracture removes more than 90% or less than 10% of an object we discard the sample and regenerate another fracture. Finally, we ensure the anomalous region of an object is visible from at least one viewpoint.
To generate a realistic scene environment, we use PBR materials [44] for floors and HDRI environment maps [56] for image-based lighting to illuminate scenes. We randomly select a pair of PBR material and HDRI environment maps from the assets to randomize the scene background. Objects are placed randomly such that no collisions occur, and each object’s rotational poses are obtained using Blender’s rigid body simulation [43]. We employed Blender 2.93 [43] with Cycles ray-tracing renderer for photo-realistic rendering. Blender 2.93 is released under the GNU General Public License (GPL, or “free software”), and the PBR and HDRI maps are released under the CC0 license.
Our framework can easily be trained with real-world manufacturing scene environments. Our model relies solely on 2D supervision, making the data collection process much easier. We also do not need precise annotation of 3D bounding boxes as they are not used during training. For annotating instance-wise anomaly labels, we can use 2D bounding boxes, which can be projected in 3D using Visual Hull [28], then used for coarse localization.
A.2 Training Details
We train our model in two stages. In the first stage, we train it with just image and feature reconstruction losses. In the second stage, we train the model end-to-end with both reconstruction and binary classification losses. All experiments are performed on a single NVIDIA A40 GPU with a batch size of 4, utilizing 28GB of GPU memory. The first stage takes 36 hours to complete, followed by an additional 24 hours for the second stage.
The ablation experiments are conducted on the same workstation with the same GPU by removing one or a few core components from the full method. Specifically, variant methods A and B take $36$ hours to train the first stage, while only taking $14$ hours to train the second stage. Regarding variant method C, it takes similar $36$ and $24$ hours for the two stages as in the full method.
We compare our method with COLMAP (BSD license), ImVoxelNet (MIT license), and DETR3D (MIT license). For the COLMAP-based approach, we use DGCNN (MIT license) as a feature extractor. In terms of training time of these baseline methods, COLMAP takes $10$ hours to train, ImVoxelNet takes $24$ hours to train, and DETR3D takes $2$ days to converge.
The standard deviation of all our experiments (including the ablations and our method) under multiple runs is less than 0.5.
A.3 Additional Qualitative Results
In Fig. 9 and Fig. 10, we present additional qualitative results on ToysAD-8K and PartsAD-15K, respectively.
Figure 9: Additional results on the unseen set of ToysAD-8K dataset. For rows 1 to 3, the anomalies are easy to spot and self-explanatory. In row 4, there are two anomalies: one with broken outer parts at the back (see view 5), and another with a tilted roof (see views 1 and 2). For row 5, one leg is broken (see view 2). In row 6, the eye is missing (see view 5).
Figure 10: Additional results on the PartsAD-15K dataset.

### v2

@@ 4.1 Datasets
We have proposed two new benchmarks for multi-object AD: ToysAD-8K and PartsAD-15K. ToysAD-8K consists of real-world objects from multiple categories that enables to evaluate our model’s ability to generalize to unseen object categories. PartsAD-15K comprises a more diverse collection of mechanical object parts with arbitrary shapes, thus being free from any class-level inductive biases (e.g., a sheep with three legs are easy to be recognized as anomaly without comparison) of common objects. Both datasets include a wide range of fine-grained anomaly instances motivated by real-world applications in inspection and quality control. Scenes are generated with diverse backgrounds, illuminations, and camera viewpoints using photo-realistic ray tracing [38]. Next, we discuss the data generation for both datasets.
@@
ToysAD-8K. We start with a subset of $1050$ shapes from the Toys4K dataset [36], covering a wide range of objects across $51$ categories. Anomalies are automatically generated for each 3D shape by applying various deformations to both the geometry and texture. These include realistic cracks and fractures using [34], applying random geometric deformations [38] like bumps, bends, and twists, as well as randomly translating, rotating, and swapping materials in different parts of the shapes. This process results in a total of $2345$ anomaly shapes.
To generate each scene, we randomly select a set of objects consisting of both the normal and anomalous variants of the same instance. Most scenes contain more normal objects than anomalies and few scenes have no anomalous instances. The resting poses for the objects are generated using rigid body simulation [4] by dropping the object to the floor. Objects are scaled and placed randomly in the scene while avoiding collisions We generated a total of $8K$ scenes. Each scene consists of $3$-$6$ objects rendered in $20$ views. For the training set, we randomly select $5K$ scenes from $39$ categories. We build two disjoint test sets. The first one (seen) contains $1K$ scenes from the seen categories but with unseen object instances. The second one (unseen) contains $2K$ scenes from the rest of the $12$ novel categories.
@@
PartsAD-15K. We use a subset of the ABC dataset [23] that consists of $4200$ shapes. Anomalies are generated following the same strategy outlined above. Additionally, for each shape, we sample geometrically close instances from the dataset and assign them as anomalies to use in the same scene. This results in high-quality anomalies that closely resemble their normal counterparts in overall geometry but differ in subtle shape variations, making them anomalous in the context of the normal ones. In total, we generate $10,203$ anomaly shapes and create $15K$ scenes, each consisting of $3$ to $12$ objects rendered from $20$ different viewpoints. The dataset is split into a $12K$ training set and the remaining $3K$ scenes (unseen shapes) as the test set.
Anomaly generation and checking. Our anomaly generation process is fully automated with several quality checks to ensure the anomalies are realistic. For example, if a part detaches from the main body during positional or rotational deformations, the anomaly is discarded and regenerated with adjusted parameters. Similarly, if removing a part makes the shape disconnected (e.g., teddy bear with a floating hand after removing an arm), we discard the anomaly and try a different part. For fracture anomalies, if a particular fracture removes more than $90\%$ or less than $10\%$ of an object, it is rejected and regenerated. We also ensure that the anomalous region of an object is visible from at least one viewpoint.
Figure 3: AD qualitative results on the unseen test categories of ToysAD-8K (top row) and the test set of PartsAD-15K (bottom row) using our proposed framework. Due to limited space, two views are shown per scene. The model prediction is shown with a yellow bounding box: a 3D box for the first example (banana) and a projected box for the others for simplicity. Our model successfully predicts the correct object in all cases shown above.
@@
Datasets Recon-Recog ImVoxelNet [32] DETR3D [43] Ours
@@
4.2 Implementation details
We use a ResNet50-FPN [25] as the 2D encoder backbone and a four-scale encoder-decoder-based 3D CNN [27] as the 3D backbone. For input, we use $M$=$5$ views, each with a resolution of $256\times 256$, though our model can accept a different $M$ during inference. During training, we use $2M$ views, divided into two sets of $M$ views. One set is used to build the neural feature volume, while the other set’s camera views are used to render the results, and vice versa.
The 3D volume $\bm{F}_{v}$ contains $96\times 96\times 16$ voxels with a voxel size of $4cm$. We sample $128$ points per ray for rendering, and the rendered features ($\hat{\Phi}_{t}$) have a spatial dimension of $32\times 32$ (i.e., $h_{f}\times w_{f}$). The ground truth segmentation mask of the input image ($I_{t\sigma}$ in Eq. 3) is only used during training. The threshold applied to the density volume $\bm{V}_{\sigma}$ is set to be $0.2$ and the DBScan algorithm is run with its default parameters.
We resize the teacher’s (DINOv2) features to the same spatial dimension for loss computation. These features are precomputed for all scenes using the publicly available weights, which are not updated during distillation. These DINOv2 features are then reduced to $d_{f}$=$128$ channel dimension using PCA before distillation. We employ three sparse voxel attention blocks, and each applies 8-headed attention with $k$=$20$. The network is first pretrained for $50$ epochs with only the reconstruction loss to build a reliable initialized 3D feature volume, then trained end-to-end with both the reconstruction and the binary classification losses for another $50$ epochs. The batch size is $4$, and we use the Adam optimizer with a learning rate $2\times 10^{-5}$. The run-time of our method is $65$ms on a single A40 GPU for a typical scene with 5 views as input.
4.3 Baseline Comparisons
We evaluate the anomaly classification results using two metrics – the area under the ROC curve (AUC) and accuracy. A prediction is considered correct if the bounding box IoU is greater than 0.5 and the anomaly classification is correct. We do not use a separate localization metric since our bounding box estimations are highly accurate. These metrics are calculated object-wise and then averaged across all test scenes.
Since no prior work exists for this task, we compare our method with two relevant competitive baselines: a reconstruction-based baseline and two multi-view 3D object detection methods in  Tab. 1. For the reconstruction-based baseline, we first employ COLMAP [33] to generate a point cloud reconstruction of the multi-object scene environment. We use default dense reconstruction parameters but utilize the provided ground truth camera matrices. Given that 5 views are insufficient, we use a total of 20 views to reconstruct the scene. We extract point clouds from the reconstructed scenes for each object, and train a Siamese-style network to obtain their pairwise similarity. We use DGCNN [42] (pretrained on ShapeNet [11]) as the point cloud feature extractor and the triplet loss [18] to supervise the network. A voting strategy is then applied to aggregate the pairwise distances of all objects in the scene and classify the objects. However, this method’s performance is highly sensitive to the reconstruction quality, resulting in poor performance on both datasets.
Next, we adapt two multi-view object detection frameworks, ImVoxelNet [32] and DETR3D [43], to locate and classify each object in the scene as either anomalous or normal. ImVoxelNet constructs a voxel representation using a 2D-3D projection similar to ours, followed by a 3D detection head [24] for the final prediction. DETR3D, a transformer-based design, uses the set prediction loss [7] for end-to-end detection without non-maximum suppression. Both models are trained on our datasets and their performance is evaluated object-wise based on the output of the classification head, similar to ours. After examining the qualitative results (see supplementary), we observe that although both methods perform well for large cracks or fractures, they struggle when intra-group comparison is necessary. This is because they tend to memorize certain anomaly types and fail to capture the cross-object relations in the scene.
Our method significantly outperforms all baselines on both datasets (see Tab. 1). This highlights the effectiveness of our architecture, which is designed to match corresponding regions across objects. Notably, our model exhibits a smaller performance drop on the unseen set, thanks to the robust 3D representation that generalizes to novel categories. Qualitative results of our method are shown in Fig. 3
@@ 4.4 Ablations and Model Analysis
In this section, we present an ablation study of our model, analyze its robustness, and evaluate its performance on real scenes. Unless stated otherwise, all experiments are conducted on the ToysAD-8K unseen set using 5 input views.
Ablation of architecture design. We evaluate the main components of our model through an ablation study and report the results in Tab. 2. All variants include the 3D feature fusion module and are optimized at least for the image reconstruction loss, which is essential for constructing scene geometry. Variant A directly maps object-centric features to their binary labels using an MLP without comparing them. In contrast, variant B applies standard attention layers to the object-centric features to learn cross-object correlations. Despite the attention layers, B only shows a slight improvement (+2.1% AUC) over A.
Table 2: Ablation Results on ToysAD-8K.
@@
We then introduce DINOv2 feature distillation (variant C), which significantly improves performance by 5.8% AUC and 11.3% accuracy. This demonstrates the importance of the part-aware 3D representation and correspondences for our task (see Fig. 6). Our final design achieves further performance gain by using sparse voxel attention, which focuses on the top-$k$ most relevant features. This leverages robust correspondences learned through DINOv2 feature distillation, effectively eliminating noisy correlations and directing attention solely to corresponding object regions.
@@
Robustness. Our method demonstrates robustness to occlusions by effectively utilizing input from multiple viewpoints. As shown in Fig. 4, our model accurately identifies the anomalous region even when parts of the object are occluded in some views. In Fig. 5 (left), we show how well our model can perform with varying numbers of input views during testing. Trained with 5 views, the model is evaluated using 1, 3, 5, 10, and 20 views. The results clearly show that our model performs well with only 3-5 views, while additional views further enhance the performance.
@@
Figure 4: Resolving occlusion and 3D ambiguity using multi-view images. The anomaly ‘sheep’ in top has a missing tail (only visible in the 2nd view due to occlusion), and the ‘hammer’ handle in bottom is bent (only apparent from the 2nd view-angle due to 3D ambiguities).
Figure 5: Impact of the number of views (left) and object count (right) on model performance.
Figure 6: Correspondences are obtained in 3D using the neural feature field $\bm{V}_{f}$ and projected onto 2D views via camera matrices for visualization. The feature field is rendered at respective viewpoints, with the first 3 PCA components mapped to different color channels.
We analyze the impact of object count on model performance (see Fig. 5 (right)) using the PartsAD-15K dataset. As the number of objects in a scene increases, occlusion and lower resolution for individual objects typically occur. However, the performance drop between the two extremes is minimal ($<$ 1.5% AUC) as shown in the figure. We also assess the model’s ability to generalize to scenes with more objects than it encountered during training. To this end, we trained our model on scenes with 3-7 objects and tested on two sets: one with 3-7 objects (AUC: 86.75) and another with 8-12 objects (AUC: 85.10). These results demonstrate our model’s adaptability to varying object counts.
In another experiment, we assess our model’s performances when varying the ratio of normal to anomalous instances during testing as shown in Fig. 8. We create five scenes (a-e) by using two geometrically similar objects ob1 and ob2, and gradually introduce more instances of ob1 (1 to 5 respectively) while maintaining a total of six objects. For example, in scene (a), ob1 appears only once, making it an anomaly. Similarly, in scene (e), ob2 appears only once, hence considered an anomaly. We utilize a consistent background across all scenes to ensure uniformity. As shown in the figure, our model correctly classifies the anomalies in each scene. We note scene (c) presents an ambiguous case, where both objects appear in equal numbers. Despite this, our model is able to separate the two groups.
Real world testing. We apply our model, trained on the synthetic dataset, to a small set of real test scenes (a tabletop with three cups, four coca-cola cans, four medicine bottles, or three bananas). Each scene is captured in an indoor environment with adequate lighting using a 3D scanning software [1]. We obtain a set of input views of the scenes with globally optimized camera intrinsic and extrinsic parameters, which are then fed into our trained model. The results are shown in Fig. 7. Our model is able to correctly distinguish the anomaly objects in all four scenes.
@@
Figure 7: Four real-world scenes are tested using our method and it can successfully detect all anomalous instances. Five views are used for testing, but only one view per scene is shown.
Limitations. Our benchmark and model have a few limitations. First, this work focuses on a specific set of anomalies common in manufacturing scenarios, potentially overlooking other types of anomalies found in different applications. Due to the high cost and difficulty of obtaining real damaged objects, our dataset primarily uses shapes of synthetic objects. Additionally, our model assumes that object instances are rigid and does not handle articulations or deformations. It also assumes objects are not neither touching nor fully occluded in the scene. Furthermore, when a scene contains mostly anomalies, the lack of ‘normal’ data points for comparison (except in cases like fractures or cracks) can make detection more challenging. Lastly, during real-world scene testing, factors such as noisy camera poses and the gap between synthetic and real environments may degrade the performance.  Fig. 9 shows some failure cases of our model.
Figure 8: We create five scenes (a-e) with two similar objects, varying their counts per scene while maintaining a constant total. Our model correctly identifies minorities as anomalies in all cases except (c), where an equal number of objects creates ambiguity. Our model still selects one group. The yellow box indicates the predicted anomaly object.
Figure 9: Failure cases. We use yellow box for our model’s prediction and red box for the ground truth. In the first example, our model incorrectly predicts an anomaly, while in the second, it correctly identifies two objects but misses the third anomaly.
@@ 5 Conclusion
In this paper, we have introduced a novel AD problem inspired by real-world applications along with two new benchmarks. The proposed task goes beyond the traditional AD setting and involves a cross study of objects in a scene from multiple camera viewpoints to identify the ‘odd-looking’ instances. We show that our model is robust to varying number of views and objects, and outperforms the baselines that do not consider cross-object correlations.